# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ Abstract
We present a modified Marrone-Treanor (MMT) model for dissociation with rate parameters derived exclusively from quasiclassical trajectory (QCT) calculations on ab initio potential energy surfaces (PESs). Analysis of the QCT dataset for reactant $\mathrm{O_{2}}$ and $\mathrm{N_{2}}$ diatoms sampled from Boltzmann internal energy distributions over a wide $(T,T_{v})$-range indicates that a modified version of the classical Marrone-Treanor two-temperature model captures the most relevant physics of shock-heated dissociating diatomic species very well. We find that simple correction factors account for non-Boltzmann depletion effects observed in direct molecular simulations (DMS) employing the same PESs. The concentration-dependent functional form proposed for these correction factors ensures that depletion effects vanish at chemical equilibrium. Based on comparisons in isothermal and adiabatic heat baths we verify that the resulting MMT model accurately reproduces all major features observed in DMS, while remaining computationally inexpensive enough for large-scale computational fluid dynamics (CFD) simulations.
@@
$a_{U}$  = fitting constant in MMT model []delimited-[]\,[-][ - ]
$\mathcal{A}_{s}$  = species $s$ chemical symbol in reaction equation []delimited-[]\,[-][ - ]
$C$  = modified Arrhenius pre-exponential factor coefficient $\,[\mathrm{cm^{3}\,s^{-1}\,mol^{-1}\,K}^{-n}]$
$\mathrm{D}$  = mixture set of diatomic species []delimited-[]\,[-][ - ]
$D_{0}$  = dissociation energy $\,[\mathrm{eV}]$
$e_{v,s}$  = species $s$ specific vibrational energy $\,[\mathrm{J\,kg^{-1}}]$
$E_{v}$  = mixture vibrational energy per unit volume $\,[\mathrm{J\,m^{-3}}]$
$f_{k}^{\mathrm{NB}}$  = rate coefficient non-Boltzmann factor []delimited-[][-][ - ]
$f_{\varepsilon}^{\mathrm{NB}}$  = vibrational energy change non-Boltzmann factor []delimited-[][-][ - ]
$j$  = rotational quantum number []delimited-[][-][ - ]
$k_{\mathrm{arr}}$  = modified Arrhenius reaction rate coefficient $\mathrm{[cm^{3}\,s^{-1}\,mol^{-1}]}$
$\mathrm{k_{B}}$  = Boltzmann constant $=1.38065\times 10^{-23}\mathrm{J}$
$k_{d}$  = effective dissociation rate coefficient $\,\mathrm{[cm^{3}\,s^{-1}\,mol^{-1}]}$
$k_{r}^{f},k_{r}^{b}$  = forward and backward rate coefficient for reaction $r$ $\,\mathrm{[cm^{3}\,s^{-1}\,mol^{-1}]}$
$K_{r}^{eq}$  = reaction $r$ equilibrium constant $\,[(\mathrm{mol/cm^{3}})^{\nu_{\mathrm{T},r}}]$
$\mathrm{M}$  = arbitrary collision partner species []delimited-[]\,[-][ - ]
$\mathrm{M}_{s}$  = species $s$ molar mass $\,\mathrm{[g\,mol^{-1}]}$
$n$  = modified Arrhenius pre-exponential factor temperature exponent []delimited-[]\,[-][ - ]
$P_{d}$  = probability of dissociation []delimited-[]\,[-][ - ]
$R$  = set of chemical reactions []delimited-[]\,[-][ - ]
$\mathcal{R}_{r}^{f},\mathcal{R}_{r}^{b},\mathcal{R}_{r}$  = forward, backward and net rate per unit volume of reaction $r$ $\,[\mathrm{mol\,s^{-1}\,cm^{-3}}]$
$\vec{q}_{v}$  = vibrational energy heat flux $\,\mathrm{[W\,m^{-2}]}$
$Q$  = approximate vibrational partition function []delimited-[]\,[-][ - ]
@@
$Q_{v}^{\mathrm{relax}}$  = translation-vibration energy relaxation source term $\,\mathrm{[W\,m^{-3}]}$
@@
$Q_{v}^{\mathrm{chem}}$  = vibrational energy-chemistry coupling source term $\,\mathrm{[W\,m^{-3}]}$
@@
$Q_{v,r,s}^{\mathrm{chem}}$  = rate of species $s$ vibrational energy change due to reaction $r$ $\,\mathrm{[W\,m^{-3}]}$
$S$  = mixture set of chemical species []delimited-[]\,[-][ - ]
$\mathcal{S}_{d}$  = statistical support for dissociation []delimited-[]\,[-][ - ]
$t$  = time $\,\mathrm{[s]}$
$T,T_{\mathrm{tr}}$  = mixture translational-rotational temperature $\,\mathrm{[K]}$
$T_{D,s}$  = species $s$ dissociation temperature $\,\mathrm{[K]}$
$T_{F}$  = pseudo-temperature in MMT model $\,\mathrm{[K]}$
$T_{v}$  = mixture vibrational temperature $\,\mathrm{[K]}$
$\vec{u},\vec{v}_{s}$  = bulk and species $s$ diffusion velocities $\,\mathrm{[m\,s^{-1}]}$
$U,U^{*}$  = pseudo-temperatures in MMT model $\,\mathrm{[K]}$
[…]
$V_{D}$  = potential energy of diatom $\,\mathrm{[eV]}$
$w_{s}$  = rate of formation of species $s$ $\,\mathrm{[kg\,s^{-1}\,m^{-3}]}$
$w_{s,r}$  = rate of formation of species $s$ in reaction $r$ $\,\mathrm{[kg\,s^{-1}\,m^{-3}]}$
$Z$  = vibrational nonequilibrium rate factor []delimited-[]\,[-][ - ]
$z_{r}$  = nonequilibrium concentration ratio of reaction $r$ []delimited-[]\,[-][ - ]
$\varepsilon_{\mathrm{int}}$  = per-particle internal energy $\,\mathrm{[eV]}$
$\varepsilon_{\mathrm{rem}}$  = per-particle remaining energy for dissociation $\,\mathrm{[eV]}$
$\varepsilon_{s}(v)$  = species $s$ per-particle quantized level $v$ vibrational energy $\,\mathrm{[eV]}$
@@
$\langle\varepsilon_{v,s}\rangle_{d}$  = per-particle vibrational energy removed for species $s$ due to dissociation $\,\mathrm{[eV]}$
$\theta_{v,s}$  = species $s$ characteristic temperature of vibration $\,\mathrm{[K]}$
$\nu_{s,r}^{\prime},\nu_{s,r}^{\prime\prime}$  = species $s$ stoichiometric coefficient before and after reaction $r$ []delimited-[]\,[-][ - ]
$\nu_{\mathrm{T},r}$  = total difference of stoichiometric coefficients in reaction $r$
$\rho_{s}$  = species $s$ density $\,\mathrm{[kg\,m^{-3}]}$
@@
$\langle\tau_{s}^{\mathrm{v}}\rangle$  = species $s$ mixture-averaged vibrational relaxation time $\,\mathrm{[s]}$
@@ 1 Introduction
There are many air dissociation kinetics models in the literature, based on differing assumptions about the dominant mechanisms in dissociating air. Most of these models assume that the translational and rotational modes of the gas are close to being in equilibrium with one another, and thus the mixture translational-rotational energy is modeled by a single temperature, $T$. The vibrational energy modes relax relatively slowly, and therefore a vibrational energy conservation equation is solved, with the vibrational energy represented by a single vibrational temperature $T_{v}$. This approach was originally conceived over 60 years ago [3, 4, 5].
@@
For a two-temperature framework, three inputs are required from a dissociation model: reaction rate coefficients, characteristic vibrational relaxation times, and the change in vibrational energy due to dissociation [6]. Most dissociation models in the literature prescribe the rate coefficients and change in vibrational energy [7, 5, 8, 9, 10] and rely on vibrational relaxation times from elsewhere, such as the fits of Millikan and White [11]. However, all of these processes are tightly coupled and an accurate model requires full consistency between all three aspects of the dissociation model.
@@
Ab-initio potential energy surfaces (PESs) have recently been developed for many of the three- and four-atom systems of interest in high-temperature air chemistry. PESs describe the force on each atom during all possible spatial configurations of the atoms. Collisions of air reactants can be simulated by integrating in time on these surfaces from an initial condition to a final molecular configuration. These quasiclassical trajectories [12] (QCTs) can be used to quantify reaction dynamics, including the role of vibrational excitation on dissociation. QCT analysis can be used to sample pre-collision states from a gas at a specified translational, rotational and vibrational temperature, and obtain reaction rate coefficients by averaging over many trajectories.
@@
QCT calculations can be used to study reaction dynamics by sampling initial conditions from fixed Boltzmann or approximately-Boltzmann distributions characterized by one or more temperatures. This technique allows for focused analysis of a particular condition, or set of conditions, although the effect of non-Boltzmann distributions is not accounted for. For example, QCT analysis was employed by Bender et al. [13] to study $\mathrm{N_{2}+N_{2}}$ dissociation, as well as nitrogen dissociation with partners $\mathrm{N}$ and $\mathrm{O_{2}}$ [14, 15]. Chaudhry et al. [16, 17, 18] have also studied oxygen dissociation with partners $\mathrm{O_{2}}$, $\mathrm{O}$, and $\mathrm{N_{2}}$. For the case of $\mathrm{O_{2}+O_{2}}$ and $\mathrm{O_{2}+O}$ interactions, multiple PESs are necessary to capture the interaction of electronically ground state reactants [19, 20, 21].
@@
Using ab initio PESs, vibrational state-resolved master equation calculations [22] and direct molecular simulations (DMS) [23] make it possible to predict the dynamics of the vibrational energy distribution in a relaxing and dissociating gas mixture. It is found that the distribution of internal energies is non-Boltzmann [24, 14, 25, 26, 27], resulting in a reduction of the dissociation rate by a factor of between 2 and 5, relative to an assumed Boltzmann distribution at $T_{v}$. Thus, non-Boltzmann effects must be included in a model for air dissociation. Recent examples of both kinetic and continuum models that include non-Boltzmann effects can be found in Refs. [28, 29], where it is shown that reasonable model assumptions could lead to more simplified two-temperature models for CFD. Separate recent DMS calculations on the Minnesota PESs have also yielded characteristic vibrational relaxation times for the major diatomic species in air [30]. Thus, with accurate PESs it is possible to compute data for all three critical components of a dissociation model. In this work, we analyze a database of QCT results. We investigate the mechanics of dissociation and identify key trends that exist for the dissociation reactions considered. We find that the Marrone-Treanor (MT) model captures these trends quite well and we propose several modifications, resulting in a new multi-temperature model for air dissociation.
@@
This paper is organized as follows. In Sec. 2 we review portions of the governing fluid equations relevant to modeling finite-rate chemistry of partially dissociated air within the context of a two-temperature $(T,T_{v})$ description and introduce the MMT model. Section 3 summarizes the main findings of our QCT database analysis. We identify the most important functional dependencies for probability of dissociation and vibrational energy change per dissociation revealed by the aggregate trajectory data. In Sec. 4 we introduce the analytical functional forms used to curve-fit the two-temperature dissociation rate coefficient and vibrational energy change per dissociation to the QCT database. In Sec. 5 we discuss the need for and implementation of correction factors to account for non-Boltzmann depletion effects during quasi-steady-state dissociation in a shock-heated gas. In Sec. 6 we propose a functional dependence on mixture composition for these non-Boltzmann factors to ensure the MMT model reproduces the correct reaction and vibrational energy removal rates at thermo-chemical equilibrium. In Sec. 7 we verify our CFD implementation of MMT in $\mathrm{N_{2}/N}$ and $\mathrm{O_{2}/O}$ mixtures against DMS reference solutions in a variety of relevant scenarios. All model parameters for thermodynamics, vibrational relaxation and dissociation chemistry are directly derived from, or fully consistent with the ab initio PESs which DMS also relies upon. Finally, in Sec. 8 we formulate the conclusions of this work. A full parametrization of the MMT model for 5-species air, including vibrational relaxation time fits derived from the ab inito PESs, as well as comparisons with DMS for full air model verification is will be discussed in an upcoming companion paper.
2 The Modified Marrone-Treanor (MMT) Model
2.1 Conservation Equations for Reacting Air

### v2

# Modified Marrone-Treanor dissociation model: formulation and benchmarking for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL

@@ Abstract
We present a modified Marrone-Treanor model for dissociation with rate parameters derived exclusively from quasiclassical trajectory calculations on ab initio potential energy surfaces. Analysis of the trajectory dataset for reactant $\mathrm{O_{2}}$ and $\mathrm{N_{2}}$ diatoms sampled from Boltzmann internal energy distributions over a wide $T,T_{\mathrm{v}}$ range indicates that a modified version of the classical Marrone-Treanor two-temperature model captures the most relevant physics of shock-heated dissociating diatomic species very well. We find that simple correction factors account for non-Boltzmann depletion effects observed in direct molecular simulations employing the same potentials. The concentration-dependent functional form proposed for these correction factors ensures that depletion effects vanish at chemical equilibrium. Based on comparisons in isothermal and adiabatic heat baths we verify that the resulting two-temperature dissociation model accurately reproduces all major features observed in the direct molecular simulations, while remaining computationally inexpensive enough for large-scale computational fluid dynamics simulations.
@@
$a_{U}$  = fitting constant in MMT model []\,[-]
$\mathcal{A}_{s}$  = species $s$ chemical symbol in reaction equation []\,[-]
$C$  = modified Arrhenius pre-exponential factor coefficient $L^{3}\,T^{-1}\,N^{-1}\,\Theta^{-n}]$
$c_{0}$  = Proportionality factor $1\,\mathrm{atm\cdot s}$
$D$  = mixture set of diatomic species []\,[-]
$D_{0,s}$  = molecular dissociation energy of species $s$ $M\,L^{2}\,T^{-2}\,N^{-1}]$
$e_{\mathrm{v},s}$  = species $s$ vibrational energy per unit mass $L^{2}\,T^{-2}]$
$E_{\mathrm{v}}$  = mixture vibrational energy per unit volume $M\,L^{-1}\,T^{-2}]$
$f_{k}^{\mathrm{NB}}$  = rate coefficient non-Boltzmann factor [][-]
$f_{\varepsilon}^{\mathrm{NB}}$  = vibrational energy change non-Boltzmann factor [][-]
$j$  = rotational quantum number [][-]
$k^{\mathrm{Arr}}$  = modified Arrhenius rate coefficient $L^{3}\,T^{-1}\,N^{-1}]$
$\mathrm{k_{B}}$  = Boltzmann constant $1.38065\times 10^{-23}\,\mathrm{J\cdot K^{-1}}$
$k_{\mathrm{diss/rec}}$  = effective dissociation/recombination rate coefficient $L^{3}\,T^{-1}\,N^{-1}]$
$k_{r}^{f/b}$  = forward / backward rate coefficient for generic reaction $r$ $(L^{-3}\,N)^{1-\nu_{r}^{f/b}}\,T^{-1}\,]$
$K_{\mathrm{c}}^{\mathrm{eq}}$  = molar concentration-based equilibrium constant $(L^{-3}\,N)^{\nu_{r}^{\mathrm{T}}}]$
$L$  = physical dimension of length
$M$  = physical dimension of mass
$\mathrm{M}$  = arbitrary collision partner species []\,[-]
$M_{s}$  = species $s$ molar mass $M\,N^{-1}]$
$m_{s,q}^{\mathrm{high/low}}$  = high/low-temperature limit vibrational relaxation time fit slope parameter for species pair $s,q$ $\Theta^{1/3}]$
$n_{s,q}^{\mathrm{high/low}}$  = high/low-temperature limit vibrational relaxation time fit offset parameter for species pair $s,q$ []\,[-]
NN  = physical dimension of amount of substance
$n$  = modified Arrhenius pre-exponential factor temperature exponent []\,[-]
$n_{\mathrm{v},s}$  = number of discrete vibrational levels of diatomic species $s$ []\,[-]
$p$  = mixture static pressure $M\,L^{-1}\,T^{-1}]$
$P_{d}$  = probability of dissociation []\,[-]
$P_{d}\,(v)$  = vibration-specific probability of dissociation []\,[-]
$R$  = set of chemical reactions []\,[-]
$r_{\mathrm{A-A}}$  = inter-atomic distance between atom pair A-A [L]\,[L]
$\mathcal{R}_{r}^{f},\mathcal{R}_{r}^{b},\mathcal{R}_{r}$  = forward, backward and net rate per unit volume of reaction $r$ $L^{-3}\,T^{-1}\,N]$
$\vec{q}_{\mathrm{v}}$  = vibrational energy heat flux $M\,T^{-3}]$
$Q_{\mathrm{v},s}$  = vibrational partition function of species $s$ []\,[-]
$\tilde{Q}_{\mathrm{v},s}$  = approximate vibrational partition function of species $s$ []\,[-]
$S$  = mixture set of chemical species []\,[-]
$\mathcal{S}_{d}$  = statistical support for dissociation []\,[-]
$t$  = time [T]\,[T]
$T$  = physical dimension of time
$T$  = mixture translational-rotational temperature $\Theta]$
$T_{\mathrm{D},s}$  = species $s$ characteristic dissociation temperature $\Theta]$
$T_{F}$  = pseudotemperature in MMT model $\Theta]$
$T_{\mathrm{r},s}$  = species $s$ rotational temperature in DMS calculation $\Theta]$
$T_{\mathrm{t}}$  = mixture translational temperature in DMS calculation $\Theta]$
$T_{\mathrm{t-r}}$  = mixture translational-rotational temperature in DMS calculation $\Theta]$
$T_{\mathrm{v},s}$  = species $s$ vibrational temperature in DMS calculation $\Theta]$
$T_{\mathrm{v}}$  = mixture vibrational temperature $\Theta]$
$\vec{u},\vec{v}_{s}$  = bulk and species $s$ diffusion velocities $M\,L^{-1}]$
$U,U^{*}$  = pseudotemperatures in MMT model $\Theta]$
$v$  = vibrational quantum number $\mathrm{[-]}$
$V_{D}$  = potential energy of diatom $M\,L^{2}\,T^{-2}]$
$w_{s}$  = rate of formation of species $s$ mass per unit volume $M\,L^{-3}\,T^{-1}]$
$w_{s,r}$  = rate of formation of species $s$ mass per unit volume in reaction $r$ $M\,L^{-3}\,T^{-1}]$
@@
$w_{\mathrm{v}}^{\mathrm{relax}}$  = vibration-translational energy relaxation source term $M\,L^{-1}\,T^{-3}]$
@@
$w_{\mathrm{v}}^{\mathrm{chem}}$  = vibrational energy-chemistry coupling source term $M\,L^{-1}\,T^{-3}]$
@@
$w_{\mathrm{v},r,s}^{\mathrm{chem}}$  = rate of species $s$ vibrational energy change due to reaction $r$ $M\,L^{-1}\,T^{-3}]$
$x_{s}$  = mole fraction of species $s$ []\,[-]
$Z$  = vibrational nonequilibrium rate factor []\,[-]
$\varepsilon_{\mathrm{int/rot/vib}}$  = per-molecule internal/rotational/vibrational energy $M\,L^{2}\,T^{-2}]$
$\varepsilon_{\mathrm{rem}}$  = per-molecule remaining energy for dissociation $M\,L^{2}\,T^{-2}]$
$\varepsilon_{s}(v)$  = species $s$ per-molecule quantized level $v$ vibrational energy $M\,L^{2}\,T^{-2}]$
@@
$\langle\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$  = per-molecule vibrational energy removed for species $s$ due to dissociation $M\,L^{2}\,T^{-2}]$
$\langle\Delta\varepsilon_{\mathrm{v},s}\rangle_{\mathrm{diss}}$  = per-molecule vibrational energy change for species $s$ due to dissociation $M\,L^{2}\,T^{-2}]$
$\zeta_{r}$  = nonequilibrium concentration ratio of reaction $r$ []\,[-]
$\Theta_{\mathrm{v},s}$  = species $s$ characteristic temperature of vibration $\Theta]$
$\Theta$  = physical dimension of temperature
$\nu_{s,r}^{f/b}$  = species $s$ stoichiometric coefficient in forward / backward sense of reaction $r$ []\,[-]
$\nu_{r}^{f/b}$  = stoichiometric coefficient sum in forward / backward sense of reaction $r$: $\nu_{r}^{f/b}=\sum_{s\in S}\{\nu_{s,r}^{f/b}\}$$\mathrm{[-]}$
$\nu_{r}^{\mathrm{T}}$  = total difference of stoichiometric coefficients in reaction $r$: $\nu_{r}^{\mathrm{T}}=\nu_{r}^{b}-\nu_{r}^{f}$$\mathrm{[-]}$
$\rho_{s}$  = species $s$ density $M\,L^{-3}]$
@@
$\langle\tau_{s}^{\mathrm{v}}\rangle$  = species $s$ mixture-averaged vibrational relaxation time [T]\,[T]
$\tau_{s,q}^{\mathrm{v}}$  = species $s$ vibrational relaxation time with collision partner species $q$ [T]\,[T]
CFD  = Computational Fluid Dynamics
DMS  = Direct Molecular Simulation
MMT  = Modified Marrone-Treanor
M&W  = Millikan and White
PES  = potential energy surface
QCT  = quasi-classical trajectory
QSS  = quasi-steady state
SHO  = simple harmonic oscillator
@@ 1 Introduction
There are many air dissociation kinetics models in the literature, based on differing assumptions about the dominant mechanisms in dissociating air. Most of these models assume that the translational and rotational modes of the gas are close to being in equilibrium with one another, and thus the mixture translational-rotational energy is modeled by a single temperature, $T$. The vibrational energy modes relax relatively slowly, and therefore a vibrational energy conservation equation is solved, with the vibrational energy represented by a single vibrational temperature $T_{\mathrm{v}}$. This approach was originally conceived over 60 years ago [3, 4, 5].
@@
For a two-temperature framework, three inputs are required from a dissociation model: reaction rate coefficients, characteristic vibrational relaxation times, and the change in vibrational energy due to dissociation [6]. Most dissociation models in the literature prescribe the rate coefficients and change in vibrational energy [7, 5, 8, 9, 10] and rely on vibrational relaxation times from elsewhere, such as the fits of Millikan and White [11] (M&W). However, all of these processes are tightly coupled and an accurate model requires full consistency between all three aspects of the dissociation model.
@@
Ab initio potential energy surfaces (PESs) have recently been developed for many of the three- and four-atom systems of interest in high-temperature air chemistry. Such PESs describe the force on each individual atom during any arbitrary spatial arrangement of the multi-atom system. Collisions of air reactants can be simulated by integrating in time on these surfaces from an initial condition to a final post-collision molecular configuration. These quasiclassical trajectories [12] (QCTs) can be used to quantify reaction dynamics, including the role of vibrational excitation on dissociation. QCT analysis can be used to sample pre-collision states from a gas at a specified translational, rotational and vibrational temperature, and obtain reaction rate coefficients by averaging over many trajectories. Calculations for the species pairings considered in this work were conducted entirely on a set of potentials generated by the University of Minnesota Computational Chemistry group. Trajectory calculations for $\mathrm{N_{2}-N_{2}}$ pairs were performed on the electronically adiabatic $\mathrm{N_{4}}(^{1}A)$ PES [13, 14, 15]. The same PES was originally used to simulate $\mathrm{N_{2}-N}$ trajectories by keeping the fourth $\mathrm{N}$ atom at effectively infinite distance from the interacting $\mathrm{N_{3}}$ sub-system. A recently released, dedicated $\mathrm{N_{3}}(^{4}A^{\prime\prime})$ PES for ground-electronic-state $\mathrm{N_{2}-N}$ collisions [16] now supersedes this approach. Trajectory calculations for $\mathrm{O_{2}-O_{2}}$ pairs were carried out on three $\mathrm{O_{4}}$ PESs by Paukku et al. [17, 18], whereas a total of 9 ground-electronic-state $\mathrm{O_{3}}$ PESs were necessary to fully simulate $\mathrm{O_{2}-O}$ trajectories [19]. For $\mathrm{N_{2}-O_{2}}$ pairs the triplet surface by Varga et al. [20] was employed.
@@
In prior works, QCT calculations on the Minnesota potentials have been used to study reaction dynamics by sampling initial conditions from fixed Boltzmann or approximately-Boltzmann distributions characterized by one or more temperatures (i.e. sweeps over $T$, $T_{\mathrm{v}}$ ranges). This approach allowed for focused analysis of a particular condition, or set of conditions, although the effect of non-Boltzmann distributions was not accounted for. For example, QCT analysis was employed by Bender et al. [15] to study $\mathrm{N_{2}+N_{2}}$ dissociation, producing estimates of thermal and multi-temperature dissociation rate coefficients, $k_{\mathrm{diss}}(T)$ and $k_{\mathrm{diss}}(T,T_{\mathrm{v}})$, as well as average vibrational energy removed per dissociation $\langle\varepsilon_{\mathrm{v,N_{2}}}\rangle_{\mathrm{diss}}$. Similar analyses were conducted for nitrogen dissociation with partners $\mathrm{N}$ and $\mathrm{O_{2}}$ [21, 22]. Chaudhry et al. [23, 24, 25] studied oxygen dissociation with partners $\mathrm{O_{2}}$, $\mathrm{O}$, and $\mathrm{N_{2}}$, producing equivalent estimates of $k_{\mathrm{diss}}(T,T_{\mathrm{v}})$ and $\langle\varepsilon_{\mathrm{v,O_{2}}}\rangle_{\mathrm{diss}}$ for these reactions. Most of these data have been collected and summarized in Ref. [25].
@@
Beyond this approach, relying on ab initio PESs in rovibrational state-resolved master equation calculations [26, 27, 28] and Direct Molecular Simulations (DMS) [29] makes it possible to predict the dynamics of the vibrational energy distribution in a dynamically relaxing and dissociating gas mixture. Over multiple DMS studies [30, 21, 31, 32, 33, 34] it has consistently been observed that the distribution of internal energies is non-Boltzmann, resulting in a reduction of the dissociation rate by a factor of $2$ to $5$ relative to an assumed Boltzmann distribution at $T_{\mathrm{v}}$. Thus, non-Boltzmann effects must be accounted for in some capacity in any practical multi-temperature model for air dissociation. Recent examples of both kinetic and continuum models that include non-Boltzmann effects can be found in Refs. [35, 36], where it is shown that reasonable model assumptions could lead to more simplified two-temperature models for Computational Fluid Dynamics (CFD). Recent DMS calculations of air dissociation in the quasi-steady-state (QSS) regime [34] in particular, have yielded estimates for non-Boltzmann rate coefficients and vibrational energy removed by dissociation for the most relevant reactions in high-temperature air. A subset of these results is further analyzed in Sec. 5 to calibrate the non-Boltzmann correction factors of the two-temperature model proposed in this paper. Furthermore, DMS calculations have yielded estimates of characteristic vibrational relaxation times for the major diatomic species in air [21, 31, 32, 37]. Thus, with accurate PESs it is possible to compute data for all three critical components of a dissociation model.
@@
In this work, we analyze a database of QCT results. We investigate the mechanics of dissociation and identify key trends that exist for the dissociation reactions considered. We find that the classical Marrone-Treanor model [5] captures these trends quite well and we propose several modifications, resulting in a new multi-temperature model for air dissociation. This paper is organized as follows. In Sec. 2 we review portions of the governing fluid equations relevant to modeling finite-rate chemistry of partially dissociated air within the context of a two-temperature $(T,T_{\mathrm{v}})$ description and introduce the Modified Marrone-Treanor (MMT) model. Section 3 summarizes the main findings of our QCT database analysis. We identify the most important functional dependencies for probability of dissociation and vibrational energy change per dissociation revealed by the aggregate trajectory data. In Sec. 4 we introduce the analytical functional forms used to curve-fit the two-temperature dissociation rate coefficient and vibrational energy change per dissociation to the QCT database. In Sec. 5 we discuss the need for and implementation of correction factors to account for non-Boltzmann depletion effects during quasi-steady-state dissociation in a shock-heated gas. In Sec. 6 we propose a functional dependence on mixture composition for these non-Boltzmann factors to ensure the MMT model will reproduce the correct reaction and vibrational energy removal rates as the mixture approaches thermo-chemical equilibrium. In Sec. 7 we benchmark our CFD implementation of the MMT model against DMS reference solutions of nonequilibrium dissociation in $\mathrm{N_{2}/N}$ and $\mathrm{O_{2}/O}$ mixtures. All model parameters for thermodynamics, vibrational relaxation and dissociation chemistry are directly derived from, or fully consistent with the ab initio PESs which DMS also relies upon. Finally, in Sec. 8 we formulate the conclusions of this work.
A full parametrization of the MMT model for 5-species air, including vibrational relaxation times derived from the ab initio PESs, as well as benchmarking against DMS results for dissociating air will be discussed in an upcoming follow-up article.
2 The Modified Marrone-Treanor model
2.1 Conservation equations for a reacting gas mixture